OpenStudy (anonymous):

K is the midpoint of LM . The coordinates of K are (3, 5) and the coordinates of L are (-5, -6), find the coordinates of M.

OpenStudy (anonymous):

(x1+x2/2) , (y1+y2/2)=\(c1,c2)

OpenStudy (anonymous):

If K is the midpoint then the distance between L and K is equal to the distance between K and M, since they give you the points L and K, use the distance formula to figure out the distance between K and M and you'll get the coordinates

OpenStudy (anonymous):

take m =(x2,y2 ) as it is .... (c1,c2)=(3,5) (x1,x2)=(-5,-6)

OpenStudy (anonymous):

it would be easy if u use mid point formula .......

OpenStudy (anonymous):

so the answer is (-5,-6)

OpenStudy (anonymous):

No that's the coordinates for L

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Distance=\[\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]